1) a) Write the Taylor Series for f(x)=5. b) Write the Taylor Series for f(x) =...

1)

a) Write the Taylor Series for f(x)=5.

b) Write the Taylor Series for f(x) = 5+4x

c) Write the Taylor Series for f(x)=5+4x+7x^2

d) Write the Taylor Series for f(x)=5+4x+7x^2 + 3x^3
Answer #1

`Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

1)

T(x)=f(0)=5

2)

T(x)=f(0)/0!+f'(0)*x/1!=5+4*x

3)

T(x)=f(0)/0!+f'(0)*x/1!+f''(0)*(X^2)/(2!)=5+4*x+7*x^2

4)

T(x)=f(0)/0!+f'(0)*x/1!+f''(0)*(X^2)/(2!)+f'''(0)*(x^3)/3!=5+4*x+7*x^2+3*x^3

Kindly revert for any queries

Thanks.